The width of a rectangle is 4 units less than the length.If the area is 45 square units, then find the dimensions of the rectangle
Answer by Boreal(15235)

length is x, width is x-4
Their product is 45
x(x-4)=45
x^2-4x-45=0
(x-9)(x+5)=0
only positive root is for length x to be 9 units. Width is 5 units.